Grade 12Physical Chemistry

Calculate the molar mass of the following : (i) H2O (ii) CO2 (iii) CH4

sudhanshu
12 Years agoGrade 12

1 Answer

Gaurav
11 Years ago
Hello Student
Molar mass of H2O = 2X Atomic mass of H + Atomic mass of O = 2x1 +16 =18
Similarly
Molar mass of CO2 = 12+2x16 = 44
and
Molar mass of CH4 = 12+4X1 = 16